Main INDEX
Monthly INDEX
PREV
NEXT
Make New Entry,
Make Followup Entry
User name R. Subedi
Log entry time 23:50:59 on December21,2008
Entry number 254342
keyword=PVDIS parasitic activities on Solid target
22:20 Beam has been off since more than 2 hours. Waiting to take parasitic run on a solid (BeO/optics ) target thinking that we could see high signal rate in the detectors. With He3 target, we used to get about 30 Hz of T8 (S1 & S2) at the current kinematic setting of right HRS (momentum 200 MeV, angle -90 deg, only Q2 and Dipole ON).
23:29 Asked for 2 uA beam (see figure 2) on BeO/optics target. Started the right HRS run 20221. Rate was only about 20 Hz of T8 (see figure 1). Took the run only for 5 minutes (5892 events). The run was not so exciting since we did not have usual high current (12 uA), and due to this small current, we were not able to see high rate.
Conclusion: we did not see any signal in the detector stack at this low rate (20 Hz). While looking at the histograms, there is no detector response, saw only pedestals (not shown here).
FIGURE 1
FIGURE 2